# SWI: Speaking with Intent in Large Language Models

### v1

@@ 7 Conclusion
In this work, we introduce Speaking with Intent (SWI) in LLMs, where the generated intent (as high-level planning) guides subsequent analysis, improving the reasoning and generation abilities. Extensive experiments across mathematical reasoning, multiple-choice QA, and text summarization benchmarks consistently show the benefits of SWI over the Baseline method that generates without explicit intent. In addition, SWI outperforms answer-trigger prompting methods like Chain-of-Thought on mathematical reasoning benchmarks. In text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations. Furthermore, human evaluations solidify the coherence, effectiveness, and interpretability of LLM-generated intent. This study opens a new avenue for enhancing LLM reasoning abilities and beyond.
Impact Statement.
As intent is a fundamental aspect of natural language processing, empowering, eliciting, and enhancing the intent understanding and generation abilities can potentially drive AI systems (including multimodal models) to the next level. Moreover, Speaking with Intent can also be applied to various domains beyond NLP, such as healthcare, law, and finance. These applications are cost-sensitive, so explicitly showing the intent of AI models will help with the transparency and interpretability of critical decision-making.
@@ Acknowledgments
Appendix A Reproducibility Statement
A.1 LLM Generation
For all experiments in § 4 and § 5, we employ LLaMA3-8B-Instruct (Dubey et al., 2024), an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters, and use the model checkpoint and tokenizer [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct provided by Hugging Face Transformers (Wolf et al., 2020). Only in the fact-checking evaluation for summaries (§ 5.2.3), we adopt the proprietary model GPT-4o-mini (Hurst et al., 2024) to decompose the generated summary and reference summary into two sets of atomic facts.
For each running session, the experiments are conducted on a single NVIDIA V100 GPU with 32GB memory. To avoid out-of-memory issues, all the models are loaded in a half-precision (float16) mode, and the generation batch size is $1$. The input sequence is not truncated since we do not want to lose the context information, but we set the maximum number of newly generated tokens as $4096$ during generation.
To guarantee reproducibility, we fixed the seeds as 42 for all random modules, set the LLM generation temperature as 0 for deterministic generation without sampling, and ran all experiments twice, obtaining reproducible generation results and evaluation scores. The source code is available on GitHub: https://github.com/YuweiYin/SWI
A.2 Experimental Cost
In the generation stage, the total computational cost is approximately 1,500 GPU hours on NVIDIA V100 clusters (about 58 days). In the evaluation stage, only the option selection process for multiple-choice QA tasks requires GPU usage (Yin & Carenini, 2025), and the overall running time is roughly 100 hours on V100 clusters. In the fact-checking experiments (§ 5.2.3), the expense for GPT-4o-mini API calls is below US\$10.
Appendix B Tasks and Datasets
To comprehensively study the effectiveness and generalizability of the proposed SWI method, we consider three different categories of tasks in our experiments: mathematical reasoning, multiple-choice QA, and text summarization. Table 10 presents the dataset sources. Please note that the URLs may be subject to change by the dataset providers.
Task Dataset URL
Mathematical Reasoning GSM8K (Cobbe et al., 2021) Link
@@ Appendix B Tasks and Datasets
AMC23 (American Mathematics Competitions) Link
AIME24 (American Invitational Math Examination) Link
AIME25 (American Invitational Math Examination) Link
Multiple-choice QA LogiQA (Liu et al., 2020) Link
CSQA (Talmor et al., 2019) Link
SIQA (Sap et al., 2019) Link
OBQA (Mihaylov et al., 2018) Link
ARC (Clark et al., 2018) Link
BBH (Suzgun et al., 2023) Link
MMLU (Hendrycks et al., 2021a) Link
MMLU-Pro (Wang et al., 2024e) Link
Text Summarization CDM (Hermann et al., 2015) Link
XSum (Narayan et al., 2018) Link
XL-Sum (Hasan et al., 2021) Link
SAMSum (Gliwa et al., 2019) Link
DialogSum (Chen et al., 2021) Link
WikiLingua (Ladhak et al., 2020) Link
@@
B.1 Mathematical Reasoning
In the mathematical reasoning task, the model is asked to solve the given math problem and present the final answer. We consider the following math datasets.
GSM8K.
Grade School Math 8K (GSM8K) (Cobbe et al., 2021) is a dataset of 8.5K high-quality linguistically diverse grade school math word problems. The dataset was created to support the task of question answering on basic mathematical problems that require multi-step reasoning.
GSM8K-P.
GSM8K-Platinum (Vendrow et al., 2025) is a revised version of the full test set of GSM8K (Cobbe et al., 2021). It revises the labels of mislabeled examples and removes any question that is determined to be poorly written (mostly due to ambiguity in the problem statement), providing a more accurate assessment of mathematical reasoning capabilities.
MATH500.
MATH500 (Lightman et al., 2024) consists of 500 questions from the MATH test set (Hendrycks et al., 2021b). The distribution of difficulty levels and subjects in MATH500 is representative of the MATH test set as a whole.
Competition.
We consider three competition-level math benchmarks: AMC23, AIME24, and AIME25, drawn from American Mathematics Competitions and American Invitational Mathematics Examination. AMC23, AIME24, and AIME25 have 40, 30, and 30 math problems, respectively. Despite the smaller scale, these math problems are much more challenging than those in GSM8K and MATH500.
B.2 Multiple-choice QA
Since many challenging benchmarks are designed as multiple-choice question-answering tasks, our SWI method is also evaluated on various reasoning-intense QA datasets, where the model is asked to select the most appropriate one from the given options to answer the question.
LogiQA.
LogiQA Liu et al. (2020) is a reading comprehension dataset that requires the model to have logical reasoning for question-answering.
CSQA.
CommonsenseQA (CSQA) (Talmor et al., 2019) examines the model on commonsense question-answering problems constructed using information from ConceptNet Speer et al. (2017).
SIQA.
SocialIQA (SIQA) (Sap et al., 2019) is a large-scale QA benchmark for commonsense reasoning about social situations, which probes emotional and social intelligence in everyday situations.
OBQA.
OpenBookQA (OBQA) (Mihaylov et al., 2018) asks the model to answer the question based on the given elementary-level science facts and broad commonsense knowledge.
ARC.
AI2 Reasoning Challenge (ARC) (Clark et al., 2018) contains grade-school science questions. It is divided into a Challenge and an Easy set, where the former contains only questions answered incorrectly by both a retrieval-based algorithm and a word co-occurrence algorithm.
BBH.
BIG-Bench Hard (BBH) (Suzgun et al., 2023) is a suite of challenging tasks filtered from BIG-Bench (Srivastava et al., 2023). Solving these problems often requires multi-step reasoning. In this work, $4$ (out of $27$) subtasks in BBH (word_sorting, object_counting, dyck_languages, and multistep_arithmetic_two) are discarded as they are not multiple-choice QA tasks.
MMLU.
MMLU (Hendrycks et al., 2021a) comprehensively measures the multitask language understanding abilities on $57$ subtasks including elementary mathematics, history, computer science, and more.
MMLU-Pro.
MMLU-Pro (Wang et al., 2024e) extends the mostly knowledge-driven MMLU benchmark by integrating more challenging, reasoning-focused questions and expanding the choice set from four to ten options.
B.3 Text Summarization
Beyond math reasoning and multiple-choice QA, we hypothesize that Speaking with Intent benefits natural language generation tasks like summarization, where the generated intent can guide the model in summarizing the source article point by point. Hence, we test the effect of SWI on the following text summarization datasets (English only).
CNN/DailyMail.
CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017) contains over 300K unique news articles written by journalists at CNN and the Daily Mail.
XSum.
Extreme summarization (XSum) (Narayan et al., 2018) is a real-world, large-scale dataset consisting of online articles from the British Broadcasting Corporation (BBC).
XL-Sum.
XL-Sum (Hasan et al., 2021) is a comprehensive and diverse dataset comprising 1.35 million professionally annotated article-summary pairs from BBC News. The dataset covers 45 languages and we only use the English subset.
SAMSum.
SAMSum (Gliwa et al., 2019) contains about 16K messenger-like conversations with summaries. Linguists were asked to create conversations similar to those they write on a daily basis, reflecting the proportion of topics of their real-life messenger conversations.
DialogSum.
DialogSum (Chen et al., 2021) is a large-scale dialogue summarization dataset, consisting of 13,460 dialogues with corresponding manually labeled summaries and topics.
WikiLingua.
WikiLingua (Ladhak et al., 2020) is a large-scale, multilingual dataset for evaluating cross-lingual abstractive summarization systems. It extracts article and summary pairs in 18 languages from wikiHow, and we only use the English subset.
Appendix C Human Evaluation Details
@@
We hire human evaluators from the cloud-sourcing platform CloudResearch [^5] https://www.cloudresearch.com/ to conduct human evaluation on the quality of the generated intent: coherence, effectiveness, and interpretability. To ensure the annotation quality, we apply several requirements to select qualified human evaluators, as shown in Table 11.
@@
For each task category, we select two datasets: GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2024) for mathematical reasoning, BBH (Suzgun et al., 2023) and MMLU (Hendrycks et al., 2021a) for multiple-choice QA, and CDM (Hermann et al., 2015; See et al., 2017) and XSum (Narayan et al., 2018) for text summarization. We randomly sample 12 instances per dataset and divide them into two batches of six. Each batch includes a dummy instance with deliberately reversed intents to ensure evaluators are actively engaged rather than randomly selecting responses. Evaluator submissions are accepted or rejected based on completion time and performance on the dummy instance.
@@
  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for solving the problem?

### v3

@@ 7 Conclusion
In this work, we introduce Speaking with Intent (SWI) in LLMs, where the generated intent (as high-level planning) guides subsequent analysis, improving the generation and reasoning abilities. Extensive experiments across text summarization, multi-task QA, and mathematical reasoning benchmarks consistently show the benefits of speaking with explicit intent over the direct generation baseline. In text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations. In addition, SWI outperforms existing LLM reasoning and planing methods and works synergistically with them. Further analysis substantiates the generalizability of SWI when applied to different settings. Moreover, human evaluations solidify the coherence, effectiveness, and interpretability of LLM-generated intent. Overall, this study opens a new avenue for enhancing LLM generation and reasoning abilities.
Ethics & Impact Statement
This work does not raise ethical issues, and we would like to mention the impact of SWI. As intent is a fundamental aspect of natural language processing, empowering, eliciting, and enhancing the intent understanding and generation abilities can potentially drive AI systems (including multimodal models) to the next level. Moreover, Speaking with Intent can also be applied to various domains beyond NLP, such as healthcare, law, and finance. These applications are cost-sensitive, so explicitly showing the intent of AI models will help with the transparency and interpretability of critical decision-making.
@@ Acknowledgments
Nous remercions le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) de son soutien.
@@
Appendix A Experiment Details
A.1 Dataset Details
All datasets used in this work are loaded from Hugging Face datasets. Table 12 lists the URL link of each dataset. Please note that the URLs may be subject to change by the dataset providers.
Task Dataset Source
Sum
CDM (Hermann et al., 2015) Link
XSum (Narayan et al., 2018) Link
XL-Sum (Hasan et al., 2021) Link
DialogSum (Chen et al., 2021) Link
WikiLingua (Ladhak et al., 2020) Link
QA
BBH (Suzgun et al., 2023) Link
MMLU (Hendrycks et al., 2021) Link
MMLU-Pro (Wang et al., 2024b) Link
Math
GSM8K (Cobbe et al., 2021) Link
@@ A.1 Dataset Details
A.2 Model Details
As mentioned in § 3.2, we mainly employ LLaMA3-8B-Instruct (Llama, 2024), an instruction-following LLM with 8 billion model parameters, for most experiments. In generalizability experiments (§ 5.3), we also explore LLMs of different sizes and types. Table 13 presents the URL link of each model and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020).
Model Source
LLaMA3-3B (Llama, 2024) Link
LLaMA3-8B (Llama, 2024) Link
LLaMA3-8B-R1 (Guo et al., 2025) Link
Table 13: Model Sources.
A.3 SWI Prompt Variants
As mentioned in § 2.2, we implement SWI in a straightforward prompting way for simplicity and reproducibility, i.e., we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts and restating the SWI requirement in the user prompt. In addition, we paraphrase the SWI prompt into three different versions (§ 5.4) to demonstrate that our SWI method maintains effectiveness irrespective of the specific prompt formulation. Here, we present the prompt variants in Table 14.
Type & Version Prompt Text
System Prompt
V0 (default)
You are a helpful assistant who speaks with intent. You are good at summarizing
documents and the summary must start with "Final Summary:"
During generation, follow all the requirements below:
1. Always explicitly state your own intent before speaking each sentence.
2. Each intent statement should explain the sentence that follows.
3. Your intent must start with the "<INTENT>" tag and end with the "</INTENT>" tag.
The content within the intent tags must begin with
"To" followed by a verb, such as "To accomplish a task."
4. At last, clearly and concisely give your final summary starting with "Final Summary:"
System Prompt
V1
You are a purposeful assistant skilled in document summarization who speaks with intent.
Your final response must begin with "Final Summary:"
While generating responses, adhere strictly to these instructions:
1. Before every sentence, clearly state your intent using an explanation.
2. Each intention should directly clarify the sentence that follows.
3. Use the tags "<INTENT>" and "</INTENT>" to wrap each intent statement.
Each statement inside the intent tags must begin with "To" and a verb,
for example, "To describe the process."
4. Conclude with a clear and concise final summary that begins with "Final Summary:"
System Prompt
V2
You are a helpful assistant who is skilled in text summarization and always communicates
with deliberate intent. Ensure your final output starts with "Final Summary:"
Comply with the following instructions during your response:
1. Begin each sentence with a description of your intent.
2. The intent must directly relate to and explain the sentence that comes after it.
3. Surround each intent with the tags "<INTENT>" and "</INTENT>". Each intent
statement enclosed by the tags should start with the word "To" and an action verb,
like "To explain the reasoning."
4. Finish with a succinct summary, introduced by "Final Summary:"
System Prompt
V3
You are a precise and helpful assistant proficient in text summarization, who always
speaks with deliberate intent. Your final response must begin with "Final Summary:"
While producing your response, follow these guidelines:
1. Before each sentence, declare your intent explicitly.
2. Ensure each intent explains the sentence that immediately follows.
3. Wrap every intent declaration with "<INTENT>" and "</INTENT>" tags.
Make sure that every intent statement within the tags begins with "To" and an action verb,
for example, "To justify the choice."
4. Conclude your response with a clearly stated final summary prefaced by "Final Summary:"
User Prompt
All Versions
Speak with intent and summarize the following document.
{{article}}
Table 14: SWI Prompt Variants.
Appendix B Human Evaluation Details
@@
We hire human evaluators from the cloud-sourcing platform CloudResearch to conduct human evaluation on the quality of the generated intent: coherence, effectiveness, and interpretability. To ensure the annotation quality, we apply several requirements to select qualified human evaluators, as shown in Table 15.
@@
For each task category, we select two datasets: CDM (Hermann et al., 2015; See et al., 2017) and XSum (Narayan et al., 2018) for text summarization, BBH (Suzgun et al., 2023) and MMLU (Hendrycks et al., 2021) for multi-task multiple-choice QA, and GSM8K (Cobbe et al., 2021) and MATH500 (Lightman et al., 2024) for mathematical reasoning. We randomly sample 12 instances per dataset and divide them into two batches of six. Each batch includes a dummy instance with deliberately reversed intents to ensure evaluators are actively engaged rather than randomly selecting responses. Evaluator submissions are accepted or rejected based on completion time and performance on the dummy instance.
@@
  • Effectiveness: Overall, do the intent statements help with the planning and reasoning for performing the task?